# IMPROVE: Impact of Mobile Phones on Remote Online Virtual Education

### v1

@@ 20 Hz - 30 Hz MP4 files with codec H264
The edBB platform facilitated optimal synchronization among all sensors, recording the captured samples in both UNIX time and local time, and sending the information to a common server for easy access. By using a single platform to control all sensors, it was possible to synchronize the activation and termination times of each sensor. This also allowed us to obtain precise information, such as the exact moment a learner presses a key and the corresponding video frame, without the need for additional processing.
@@
In addition to using these platforms, a log recording tool called LOGGE [22] was designed, with the primary objective of labeling the mobile phone events of Group 1 and Group 2, which consisted of learner interactions with their mobile phones or when their phones rang. Additionally, this tool was responsible for collecting information related to the monitored learner (e.g., use of glasses/contact lenses, cardiac conditions, handedness, etc.) and enhancing the edX log traces with information related to the activities the learner engages in during the learning session (start and finish times of each task, error logs during the session, etc.).
@@
Finally, a web-based system called M2LADS [22] was designed to visualize the information captured from each monitored learner. This system allowed for the following: i) Validate that the information was correctly monitored and synchronized: visualizing the information from the captured signals (webcam videos, monitor video, heart rate, attention level, etc.) and ensuring their correct synchronization. ii) Statistical analysis of the data: comparing learner performance values (pretest score vs. posttest score), analyzing signals based on activities (identifying which activities had higher attention levels or heart rates, etc.), and correlating biometric signals. The overall architecture, including the platforms and tools used to gather IMPROVE database, is summarized in Fig. 3.
@@
The videos from the RealSense RGB camera were processed to obtain the learner’s facial position, facial landmarks and 3D head posture for each frame. Three state-of-the-art methods based on Convolutional Neural Networks (CNNs) were utilized:
@@
  • Facial Detection: The facial detection module employed MediaPipe’s BlazeFace[46] model (full range) for 2D facial image detection. This module was based on CNNs specifically designed to significantly reduce computational cost and inference time without sacrificing accuracy. BlazeFace integrates a lightweight feature extraction architecture inspired by MobileNetV1/V2[47], leveraging its efficiency and speed. Specifically, the model’s architecture utilizes advanced convolution techniques, similar to those employed in networks like CenterNet[48], with a customized encoder. The detector was trained on large, private datasets that include geographically diverse data based on facial detection.
@@
  • Head Pose Detection: A real-time head pose detector based on the WHENet[49] architecture was employed. The Euler angles —vertical (pitch), horizontal (yaw), and longitudinal (roll)— were estimated through a 2D image of the head, allowing the 3D orientation of the head to be deduced from 2D facial images. This architecture was based on fully connected CNNs for angle classification. Specifically, EfficientNet[50] was utilized as the backbone for feature extraction. Several layers comprised this architecture, starting with convolutional layers that capture spatial features of the image. These convolutional layers were followed by pooling layers. A series of fully connected dense layers were used to interpret the extracted features and perform the classification and regression of the Euler angles. In the classification process, each of the angles (yaw, pitch, and roll) was classified into bins, while regression was used to accurately predict the Euler angle values within those bins. The head pose detector was trained on the 300W-LP[51] and CMU[52].
@@
  • Facial Landmark Detection: The MediaPipe library was used, which estimates 478 3D facial landmarks in real-time from facial images. This enables the extraction of relevant information from facial areas such as the eyes, forehead, mouth, and more. The model is based on a CNN, similar to MobileNetV2[47], but with custom blocks designed for real-time performance. The model was evaluated using a private dataset consisting of 1700 examples uniformly distributed across 17 geographical subregions. The mean error was 2.62, with similar error rates observed for each subregion. Consequently, each video is accompanied by the 478 3D facial landmarks for each frame, along with detailed information about the estimation of each landmark, including its visibility and presence in the image.
@@
The data obtained from the eye tracker was processed to extract relevant features such as fixations and saccades. On average, Tobii Pro Fusion captures a sample with gaze coordinates every 8.3 ms. These samples can be processed for the entire video to obtain relevant visual attention features, such as fixations and saccades. For this purpose, one of the filters integrated into the Tobii Pro Lab software was selected, specifically the I-VT (Fixation) filter. We selected this filter because it is suitable for our study, where head movements are minimal, and the predominant eye movements are fixations and saccades. This filter assigns fixations or saccades to gaze points based on their speed. If the speed is below 30 degrees per second, the points are assigned to the same fixation. Some characteristics of this filter include: i) Noise reduction: A moving average is applied every 3 samples. ii) Velocity calculation: Velocity is calculated in 20 ms windows. iii) Minimum fixation duration: Fixations shorter than 60 ms are discarded. iv) Merging adjacent fixations: Adjacent fixations are merged when two fixations occur with a time interval of 75 ms or less, and the angular distance between them is 0.5 degrees or less.
@@
Figure 4: Example of filtered signals in IMPROVE: The graph on the left shows the raw heart rate signal and the same signal filtered using a moving average filter (with a 15-second window), while the graph on the right presents the raw attention signal obtained from the EEG and filtered using a median filter (with a 5-second window). The original signals are represented by a solid blue line, while the filtered signals are shown with a solid yellow line.
@@
The data obtained from the EEG headband were filtered to mitigate the impact of potential errors, such as signal fluctuations caused by sudden movements of the participant, ensuring more accurate recordings of the EEG signals. The signals were processed using a median filter with a window size of 5, effectively reducing noise caused by artifacts, such as participant movements, eye blinks, and other factors, without distorting the peaks or extreme values in the signal. Additionally, interpolation was applied for data losses shorter than 5 seconds, while longer data gaps were marked as missing information. Both the raw and filtered data are available in the database.
@@
The heart rate data was also filtered to eliminate minor fluctuations and smooth the signal. Specifically, a moving average filter with a 15-second window was applied, as the pulse signal in this context is expected to change slowly, with variations between beats typically not being abrupt at the sampling rate of smartwatches. In this case, only one user experienced a data loss of 60 seconds, which was marked as missing information. The signal obtained from the smartwatch’s accelerometer was filtered using a fourth-order Butterworth low-pass filter with a cutoff frequency of 15 Hz. This was done to eliminate high-frequency noise, such as vibrations or unwanted movements, while preserving the rapid and relevant movements during the learner’s interactions with the mouse. For the gyroscope signal, a similar Butterworth low-pass filter was applied, but with a cutoff frequency of 10 Hz, which allowed for the capture of rapid wrist movements while eliminating potential high-frequency artifacts. Both the raw and filtered data are available in the database. Fig. 4 presents examples of filtered signals from the IMPROVE database.
@@
The IMPROVE dataset is available through Science Data Bank[53] (https://doi.org/10.57760/sciencedb.14565) and GitHub[54] (https://github.com/BiDAlab/IMPROVE). The database consists of 2.83 terabytes of information, including biometric data, physiological signals, videos, supervisor-labeled events related to mobile phone usage, and more. This information has been collected from the edBB and edX platforms, as well as from the LOGGE tool, and also includes processed signals. The dataset is organized into three main folders, each corresponding to a different group based on mobile phone usage (see Subsection Protocol for more details). Within each of these folders, there are 40 main subfolders, each labeled with the learner’s ID, containing the monitored data for each learner during the learning session. Each of these folders is further divided into four categories: ’edBB’, ’edBB Processed Data’, ’edX,’ and ’LOGGE,’ corresponding to the information collected from each of the platforms and tools. Tables summarizing the dataset structure and file descriptions—Table 6, Table 7, and Table 8—are provided in the Appendix section for reference. The processed signals and videos obtained through the edBB platform are stored in the ’edBB Processed Data’ folder and are referenced in Table 6. Table 7 lists the files and descriptions from the edX platform and the LOGGE tool. Additionally, all files from the edBB platform and their descriptions are summarized in Table 8.
Technical Validation
For the technical validation of the data, we present a se-supervised data curation method and $3$ different statistical analysis performed using the data over the different groups monitorized.
i) Semi-supervised Event Data Curation:
A semi-supervised method was employed for the data curation of the mobile phone usage events. To this end, we used the eye tracker data and head pose estimation through video (see Fig. 5) for an automatic detection of mobile phone usage events. This automatically annotated events were validated with the events labeled by the human supervisor, who manually labeled the start and end of each event.
Figure 5: Diagram of the proposed semi-supervised method for curating mobile phone usage event data.
    1. 1.
    2. 2.
    3. 3.
      A temporal window $W_{l}$ was labeled as an event when the $(\mu_{W_{l}})$ exceeds a predefined threshold, which was set based on a significant deviation from the global mean ($\mu$) of the Euler angles calculated over all sessions. The following equation shows the threshold: $\left|\mu_{W_{l}}^{i}-\mu^{i}\right|>n\sigma^{i}$, where $i$ represents the angle (yaw, pitch, or roll) and $n$ is the threshold factor that defines the sensitivity of the event detection, determining how many standard deviations away from the mean a particular value must be to be considered an event.
    4. 4.
      All overlapping events were merged.
    Figure 6 shows the detection of abnormal head pose events applied in module c), for two learners from the IMPROVE dataset.
Figure 6: Example of the abnormal head pose detector for two students. The graphs show the Euler angles obtained during the session, along with the average for each angle in a window $W_{l}=4s$. The set thresholds for each angle are also shown, and when these are exceeded, they are labeled as events. The detector exhibits a high overlap with the manually labeled events. The images show the students’ pose while using a mobile phone. Yaw is represented by the solid blue line, and roll by the solid red line. The temporal window average is indicated by the dashed yellow line. Predicted event labels are framed in red, while manually labeled events are framed in green. Thresholds are represented by dotted lines with the corresponding color for each angle.
Figure 7: Average event detection sensitivity obtained by the detect abnormal head pose approach to detect mobile phone usage across all users (left). Average number of detected events by the detect abnormal head pose approach across all users within the window $W_{s}\pm\Delta t$ (right). The x and y axes in both heatmaps represent the tested parameters $W_{l}$ and $n$, respectively, which were varied to analyze their effect on the detection performance.
Mean (Std)
Sensors Variable (Unit) Group 1 Group 2 Group 3 Overall
Accx (m/s2) 0.98 (0.48) 0.81 (0.58) 0.75 (0.50) 0.85 (0.52)
Accy (m/s2) 5.05 (1.39) 4.19 (1.54) 4.82 (1.39) 4.69 (1.44)
Accz (m/s2) 7.51 (0.89) 8.46 (0.94) 8.13 (0.95) 8.03 (0.92)
Gyrox (°/s) 0.07 (0.03) 0.04 (0.03) 0.04 (0.02) 0.05 (0.03)
Gyroy (°/s) 0.04 (0.01) 0.02 (0.01) 0.02 (0.01) 0.03 (0.01)
Gyroz (°/s) 0.04 (0.02) 0.03 (0.02) 0.03 (0.01) 0.03 (0.02)
Meditation (0-100) 56.39 (4.46) 54.25 (5.48) 54.58 (5.79) 55.07 (5.24)
Alpha (dB) 6.38 (0.62) 5.47 (0.61) 6.23 (0.61) 6.03 (0.61)
Beta (dB) 3.32 (0.62) 3.28 (0.58) 3.67 (0.58) 3.42 (0.59)
Delta (dB) 1.70 (0.97) 1.59 (1.00) 1.59 (0.95) 1.63 (0.97)
Gamma (dB) 1.89 (0.90) 1.97 (0.82) 2.06 (0.75) 1.97 (0.82)
Theta (dB) 10.02 (0.69) 9.21 (0.71) 9.84 (0.71) 9.69 (0.70)
Figure 8: Comparison of heart rate measurements between the Huawei 2 and Fitbit Sense for two different students. The solid blue line represents the Huawei 2, while the orange dashed line corresponds to the Fitbit Sense.
ii) Statistical Distributions of Physiological and Biometric Signals:
The mean and standard deviation of each biometric and physiological signal were calculated to ensure that no general errors were present and that the values fell within expected ranges for a similar environment (see Table 4). Additionally, these results were compared with the mEBAL2 dataset [21], which was obtained in a similar e-learning environment and utilized the same EEG band. The values obtained for the various physiological and biometric signals are found to be within the expected ranges for learners in an e-learning context. The average heart rate across the three groups ranged from 78.26 to 82.85 bpm, consistent with values reported in previous studies on students during virtual learning sessions [26], [14], and with the normal human resting heart rate, typically between 60-100 bpm [56].
The inertial sensors (accelerometer and gyroscope) in the Huawei Watch 2 smartwatch, worn on the hand used to operate the mouse, provided the following observations. The average acceleration in the X and Y axes reflects low values, which is expected during a learning session where most of the time is characterized by either no movement or minor mouse movements. In contrast, the higher values on the Z axis are primarily attributed to the influence of gravitational force. Similar results are observed in the gyroscope, with mean values close to zero across all axes, reflecting the same reasons.
iii) Heart Rate Data Interoperability between Devices:
The heart rate data from both smartwatches (Huawei Watch 2 and FitBit Sense) were compared to ensure consistent measurements, as similar values would indicate accurate data recording. The heart rate sampling frequency is different between the two devices, with data being captured by the Huawei Watch 2 at 1 Hz, while the FitBit Sense operates at a lower frequency of 0.2 Hz. A difference of 3.70 bpm was found between the two signals (see Fig. 8), which falls within the reasonable range for this type of device, typically around ±$\pm 3\text{-}5$ bpm [57]. Additionally, it should be noted that the smartwatches were worn on different wrists, with the dominant hand experiencing more movement, potentially affecting heart rate estimation. This, along with the variation in sampling frequencies, may have contributed to the observed differences.
Mean (Std) Statistical Tests
Sensors Variable (Unit) No Phone Phone All Sessions t p-value
Accx (m/s2) 0.74 (0.48) 2.52 (0.99) 0.98 (0.48) -10.82 0.0000
Accy (m/s2) 4.58 (1.55) 8.05 (1.03) 5.05 (1.39) -12.15 0.0000
Accz (m/s2) 8.17 (1.00) 3.31 (1.25) 7.51 (0.89) 16.80 0.0000
Gyrox (°/s) 0.05 (0.02) 0.22 (0.09) 0.07 (0.03) -11.25 0.0000
Gyroy (°/s) 0.03 (0.01) 0.09 (0.04) 0.04 (0.01) -10.50 0.0000
Gyroz (°/s) 0.04 (0.01) 0.09 (0.04) 0.04 (0.02) -9.70 0.0000
Meditation (0-100) 56.33 (4.4) 56.80 (6.51) 56.39 (4.46) -0.59 0.5626
Alpha (dB) 6.31 (1.72) 6.85 (1.78) 6.38 (1.69) -3.57 0.0011
Beta (dB) 3.30 (0.95) 3.45 (0.98) 3.32 (0.94) -2.00 0.0541
Delta (dB) 1.69 (0.43) 1.80 (0.49) 1.70 (0.43) -2.35 0.0250
Gamma (dB) 1.86 (0.80) 2.03 (0.72) 1.89 (0.78) -2.36 0.0243
Theta (dB) 9.99 (2.36) 10.18 (2.13) 10.02 (2.31) -0.88 0.3839
iv) Statistical Validation of Signal Changes Due to Mobile Phone Usage:
Notably, the accelerometer and gyroscope signals exhibit highly significant differences between the phone and no-phone conditions, as expected due to the movement involved when using a mobile phone, which contrasts with typical movements during an online session. Heart rate also shows a significant difference, indicating that the heart rate is affected by mobile phone use. Most EEG signals, including attention, also show significant differences. "However, the Meditation, Beta, and Theta signals do not exhibit statistically significant differences with respect to the mean. Theta waves are related to relaxation and wakeful states, so it is unsurprising that the Meditation signal shows no significant differences, suggesting that relaxation or meditation is not notably affected by mobile phone usage. However, further research is required to confirm this.
Figure 9 presents boxplots for attention and Theta waves across two experimental conditions: "Without Using Phone" and "Using Phone." Attention levels tend to be lower when the mobile phone is used, while the range of values is wider when the phone is not in use. In contrast, Theta waves show a slight increase during mobile phone usage.
iv) Verification of Signals and Video Recordings with M2LADS
M2LADS [22] was used to validate the signals and video recordings captured during the session. Through the system’s visualization and the synchronization it provides, all data were checked, and no corrupt files were reported by the system. The results of the validation process using M2LADS were as follows:
  • EEG Band: 35% of the learners experienced some data loss, with 23.3% of users losing 1 minute of data and 14.16% losing up to 5 minutes of data. These losses were primarily caused by abrupt head movements, which dislodged the EEG band, occurring more frequently among female participants. Supervisors closely monitored the sessions to prevent such issues.
  • Depth Video: One depth video file was found to be corrupt for the user with ID: 202303281.
  • Left NIR Camera: A corrupt file was identified for the left NIR camera in the session for the user with ID: 202305233.
  • Frontal RGB Camera: It was reported that an RGB camera video (ID: 202307122) continued recording a few minutes after the session had ended, instead of stopping at the session’s conclusion.
  • Fitbit Sense: The Fitbit Sense data were lost for three students (IDs: 202303272, 202303273, and 202303274).
Figure 9: Boxplot comparing mobile phone usage and non-usage conditions for attention levels (left) and theta wave activity (right).
Usage Notes
Potential applications. The IMPROVE dataset offers a unique opportunity for researchers across various disciplines. Its multimodal data, including physiological signals, video recordings, learning analytics, and learner feedback, provide numerous potential applications. The primary goal of the IMPROVE dataset is to assess the effects of mobile phone usage on learners during online learning sessions. IMPROVE serves as a valuable resource to explore how interaction with mobile phones affects cognitive performance and learner well-being.
Machine learning and signal processing researchers can leverage this dataset to develop and refine models that detect the impact of mobile phone usage on learning performance. The combination of different data sources, such as EEG bands, inertial signals, gaze tracking, heart rate monitors, and video from cameras, enables the exploration of complex patterns that could lead to enhanced algorithms for estimating attention, cognitive load, and the distractions caused by mobile phone usage.
This dataset is also valuable for educational researchers studying the interaction between physiological responses and online learning environments. By analyzing both biometric signals and academic performance data, researchers can investigate the effects of mobile phone usage on cognitive performance, attention span, and learning outcomes.
Similar to the edBB dataset [14, 1], IMPROVE opens up challenges to the video processing community, including tasks related to RGB, NIR, and depth camera data, such as pulse estimation, blink detection, user pose estimation, and detection of abnormal learner behavior, among others.
Moreover, researchers in Human-Computer Interaction (HCI) could utilize the IMPROVE dataset to develop adaptive learning systems that respond to the physiological states of learners [1, 22].
Psychology and neuroscience communities can also benefit from this dataset by investigating mobile phone use’s cognitive and emotional effects in educational settings. EEG signals, particularly those related to attention and relaxation (e.g., Theta and Beta waves), can be analyzed to explore how cognitive processes are impacted during online sessions involving varying levels of mobile phone interaction.
In summary, the IMPROVE dataset offers extensive potential for applications in educational technology, machine learning, psychology, and human-computer interaction, making it a valuable resource for interdisciplinary research into the effects of mobile phone usage in digital learning environments.
Suggestions: It is highly recommended to use the filtered EEG band files or apply filters using programs like Matlab or Python. In Python, the SciPy library (https://scipy.org) provides algorithms for signal processing, such as signal filtering. The EEG band is often affected by artifacts due to movement or eyeblinks, so it is recommended to apply a moving average or median filter to improve the quality of the signals [8].
The Pandas library (https://pandas.pydata.org/) proved very useful for manipulating and analyzing large datasets, while NumPy (https://numpy.org) was used for processing any type of signal or video. For training deep learning models or neural networks, TensorFlow (https://www.tensorflow.org/) and PyTorch (https://pytorch.org/) are two of the most powerful and flexible platforms for this purpose, and were essential for running the models applied in video processing.
For video processing, the Dlib library (www.dlib.net), which includes pre-trained models for facial detection, facial landmark estimation, and more, was used. This library integrates well with OpenCV (https://opencv.org/), which facilitates video reading, filtering, distortion correction, and also provides pre-trained models. This combination was used to process the videos, alongside MediaPipe (https://github.com/google-ai-edge/mediapipe?tab=readme-ov-file), a library specializing in fast models for face detection, facial landmarks, image segmentation, and more. MediaPipe was employed to extract the facial bounding boxes and landmarks available in the IMPROVE dataset. The use of RetinaFace (https://github.com/serengil/retinaface) is also recommended for facial detection with facial landmarks.
For head pose estimation, WHENet (www.github.com/Ascend-Research/HeadPoseEstimation-WHENet) is a fast and efficient solution, which was used to obtain the Euler angle labels in IMPROVE. Another efficient option for head pose estimation is RealHePoNet (https://github.com/rafabs97/headpose_final).
For processing eye-tracker data, the official software Tobii Pro Lab (https://www.tobii.com/products/software/behavior-research-software/tobii-pro-lab) is recommended, which was used in IMPROVE. Another interesting solution is PsychoPy (https://www.psychopy.org/), a free tool offering support for the creation and execution of experiments involving eye tracking.
Data Access: The IMPROVE dataset is available exclusively for research purposes and must not be used or included in any commercial applications in any form (e.g., original files, encrypted files, files containing extracted features, etc.). Researchers wishing to access the dataset can do so through two available methods. The license agreement can be downloaded from GitHub (https://github.com/BiDAlab/IMPROVE) and a signed, scanned copy should be sent via email to atvs@uam.es, following the instructions provided on the GitHub repository. Alternatively, access can be requested through the Science Data Bank (https://doi.org/10.57760/sciencedb.14565), where the necessary steps for obtaining the data are outlined on the platform.
Code Availability
The code for filtering physiological and biometric signals, as well as detecting pose events, is available on GitHub (https://github.com/BiDAlab/IMPROVE). MATLAB was used for signal filtering, while Python was utilized for pose event detection. All necessary packages are listed in the requirements.txt file. These available codes aim to assist the community in processing the IMPROVE dataset for various research and analysis purposes.
Acknowledgements
This work has been supported by projects: HumanCAIC (TED2021-131787B-I00 MICINN), BIO-PROCTORING (GNOSS Program, Agreement Ministerio de Defensa-UAM-FUAM dated 29-03-2022), SNOLA (RED2022-134284-T), e-Madrid-CM (S2018/TCS4307). Research partially funded by the Autonomous Community of Madrid. Roberto Daza is supported by a FPI fellowship from MINECO/FEDER. A. Morales is supported by the Madrid Government (Comunidad de Madrid-Spain) under the Multiannual Agreement with Universidad Autónoma de Madrid in the line of Excellence for the University Teaching Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation).
Author Contributions Statement
Competing interests
The authors declare no competing interests.
Appendix: Comprehensive Overview of Data Collected and Processed in the edBB, LOGGE, and edX Platforms
@@ Acknowledgements
Table 6: Description of Processed edBB Files, Contained in the edBB Processed Data Folder.
@@
Table 7: Description of Files Collected by the edX Platform and LOGGE Tool, Contained in the edX and LOGGE Folders.

### v3

@@ 20 Hz - 30 Hz MP4 files with codec H264
The edBB platform facilitated optimal synchronization across all sensors, recording the captured samples in both UNIX time and local time, and sending the information to a common server for easy access. By using a single platform to control all sensors, it was possible to synchronize the activation and termination times of each sensor. This configuration also enabled the acquisition of precise information, such as the exact moment a learner pressed a key and the corresponding video frame, without requiring additional post-processing.
@@
In addition to using these platforms, a logging tool called LOGGE [23] was developed with the primary goal of labeling the mobile phone events for Group 1 and Group 2, which consisted of learner interactions with their mobile phones or when their phones rang. Additionally, this tool collected learner-related information (e.g., use of glasses/contact lenses, cardiac conditions, handedness, etc.) and enhanced the edX log traces with details about the activities performed during the learning sessions, such as task start and end times, error logs, and more.
@@
Finally, a web-based system called M2LADS [23, 48] was designed to visualize the information captured from each monitored learner. This system enabled the following: i) Validation of correct monitoring and synchronization: visualizing the information from the captured signals (webcam videos, screen recordings, heart rate, attention level, etc.) and ensuring their correct synchronization. ii) Statistical analysis of the data: comparing learner performance (pretest vs. posttest scores), analyzing signals across activities (e.g., identifying which activities had higher attention levels or heart rates), and correlating biometric signals. The overall architecture, including the platforms and tools used to gather IMPROVE dataset, is summarized in Fig. 3.
@@
The videos from the RealSense RGB camera were processed to extract the learner’s facial position, facial landmarks, and 3D head pose for each frame. Three state-of-the-art methods based on Convolutional Neural Networks (CNNs) were employed:
@@
  • Facial Detection: The facial detection module employed MediaPipe’s BlazeFace[49] model (full-range version) for 2D facial image detection. This module is based on CNNs specifically designed to significantly reduce computational cost and inference time without compromising accuracy. BlazeFace integrates a lightweight feature extraction architecture inspired by MobileNetV1/V2[50], leveraging their efficiency and speed. Specifically, the model’s architecture applies advanced convolution techniques, similar to those employed in networks like CenterNet[51], with a customized encoder. The detector was trained on large private datasets containing geographically diverse facial detection samples.
@@
  • Head Pose Detection: A real-time head pose detector based on the WHENet[52] architecture was employed. The Euler angles —pitch (vertical), yaw (horizontal), and roll (longitudinal)— were estimated from 2D head images, allowing the 3D orientation of the head to be deduced from 2D facial images. This architecture is based on fully connected CNNs for angle classification. Specifically, EfficientNet[53] was used as the backbone for feature extraction. The architecture consists of several layers, starting with convolutional layers that capture spatial features of the image. These convolutional layers were followed by pooling layers. A series of fully connected dense layers interprets the extracted features and performs the classification and regression of the Euler angles. In the classification process, each of the angles (yaw, pitch, and roll) was classified into bins, while regression was used to accurately predict the Euler angle values within those bins. The head pose detector was trained on the 300W-LP[54] and CMU[55].
@@
  • Facial Landmark Detection: The MediaPipe library was employed to estimate 478 3D facial landmarks in real time from facial images. This enables the extraction of relevant information from facial regions such as the eyes, forehead, mouth, and more. The model is based on a CNN, similar to MobileNetV2[50], but with custom blocks designed for real-time performance. The model was evaluated on a private dataset comprising 1,700 examples uniformly distributed across 17 geographical subregions. It achieved a normalized mean error of 2.62%, calculated relative to the inter-ocular distance, with similar error rates observed for each subregion. Consequently, each video is accompanied by the 478 3D facial landmarks for every frame, along with detailed information about the estimation of each landmark, including its visibility and presence in the image.
Eye Tracker Signal Processing.
@@
The data obtained from the eye tracker was processed to extract relevant features such as fixations and saccades. On average, Tobii Pro Fusion captures one sample with gaze coordinates every 8.3 ms. These samples can be processed over the entire video duration to obtain relevant visual attention features, such as fixations and saccades. For this purpose, one of the built-in filters provided by the Tobii Pro Lab software was selected, specifically the I-VT (Fixation) filter. We selected this filter because it is suitable for our study, where head movements are minimal and the predominant eye movements are fixations and saccades. The filter classifies gaze points as either fixations or saccades based on their velocity: points with speeds below 30 degrees per second are grouped into a single fixation. Some characteristics of this filter include: i) Noise reduction: A moving average is applied every 3 samples. ii) Velocity calculation: Velocity is calculated in 20 ms windows. iii) Minimum fixation duration: Fixations shorter than 60 ms are discarded. iv) Merging adjacent fixations: Adjacent fixations are merged if the interval between them is 75 ms or less and the angular distance is 0.5 degrees or less.
@@
Figure 4: Example of filtered signals in IMPROVE: The left graph displays the raw heart rate signal and the same signal filtered using a moving average filter (with a 15-second window). The right graph shows the raw attention signal derived from the EEG and filtered using a median filter (with a 5-second window). In both graphs, the original signals are represented by a solid blue line, while the filtered signals are shown with a solid yellow line.
EEG Signal Processing.
@@
The data obtained from the EEG headband were filtered to mitigate potential errors, such as signal fluctuations caused by sudden participant movements, ensuring more accurate EEG recordings. The signals were processed using a median filter with a window size of 5, effectively reducing noise from artifacts, such as participant movements, eyeblinks, and other sources, without distorting the peaks or extreme values in the signal. Additionally, interpolation was applied to data losses shorter than 5 seconds, while longer data gaps were marked as missing. Both the raw and filtered signals are available in the dataset.
Smartwatch Signal Processing.
@@
The heart rate data were also filtered to remove minor fluctuations and smooth the signal. Specifically, a moving average filter with a 15-second window was applied, as the pulse signal in this context is expected to change slowly, with beat-to-beat variations typically not being abrupt at the sampling rate of smartwatches. The signal obtained from the smartwatch’s accelerometer was filtered using a fourth-order Butterworth low-pass filter with a cutoff frequency of 15 Hz. This approach was used to eliminate high-frequency noise, such as vibrations or unintentional movements, while preserving the rapid and relevant movements during the learner’s interactions with the mouse. For the gyroscope signal, a similar Butterworth low-pass filter was applied, but with a cutoff frequency of 10 Hz, which allowed for the capture of rapid wrist movements while eliminating potential high-frequency artifacts. Both the raw and filtered data are available in the dataset. Fig. 4 presents examples of filtered signals from the IMPROVE dataset.
@@
The IMPROVE dataset is available from the Science Data Bank[56] (https://doi.org/10.57760/sciencedb.14565) and GitHub (https://github.com/BiDAlab/IMPROVE). The dataset comprises 2.83 terabytes of information, including biometric data, physiological signals, videos, supervisor-labeled events related to mobile phone usage, and more. This data was collected from the edBB and edX platforms and the LOGGE tool, and includes both raw and processed signals. The dataset is organized into three main folders, each corresponding to a different group based on mobile phone usage (see Subsection Protocol for more details). Within each of these folders, there are 40 main subfolders, each labeled with the learner’s ID, containing the monitored data for each learner during the learning session. Each of these subfolders is further divided into four categories: ’edBB’, ’edBB Processed Data’, ’edX,’ and ’LOGGE,’ corresponding to the information collected from each platform and tool. The structure and contents of the dataset are summarised in the following tables: Table 4, Table 5, and Table 6. The processed signals and videos obtained through the edBB platform are stored in the ’edBB Processed Data’ folder and are described in Table 4. Table 5 lists the files and descriptions from the edX platform and the LOGGE tool. Additionally, all files from the edBB platform and their descriptions are summarized in Table 6.
@@
Table 4: Descriptions of the processed edBB files included in the edBB Processed Data folder.
@@
Table 5: Description of files collected by the edX platform and LOGGE tool, included in the edX and LOGGE folders.